Graph the line that satisfies each set of conditions...
(passes through origin, parallel to graph of x+y=10

Answer by nerdybill(2709)

(passes through origin, parallel to graph of x+y=10
.
rearranging:
x+y=10
into the "slope-intercept" form:
y = (-1)x + 10
we see that the slope is -1
.
recapping, we would like a line that crosses:
(0,0) with a slope of -1
.
Plug the above into the "point slope" form and solve:
y – y1 = m(x – x1)
y – 0 = (-1)(x – 0)
y = (-1)(x)
y = -x (this is what they're looking for)